OpenStudy (anonymous):

Identify the property shown. 6j + (2k + 3) = (6j + 2k) + 3

OpenStudy (anonymous):

I'll give you a hint. By the commutative property, the following expressions are equivalent: a x b = b x a, given a, b are real numbers that are being multiplied. Distributive property is: a(b + c) = ab + bc, for real numbers a, b, c. Associative property is: a + (b + c) = (a + b) + c, for real numbers a, b, c being added. Which one do you think applies to your question? @kathy0514

OpenStudy (anonymous):

so it would be associative property?

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

oh ok thank u

OpenStudy (anonymous):

feel better?

OpenStudy (anonymous):

lol ur mean :(